# The graphene-water interface is acidic

### v1

@@ Conclusion
Our findings demonstrate a clear preference for the hydronium ion at the graphene-water interface under various confinement regimes, rendering the graphene-water interface acidic. Enthalpic forces drive the hydronium ion to the interface, maintaining the hydrogen bonding network and enabling energetically favorable interactions with graphene. Conversely, a subtle balance of entropic and enthalpic contributions generally pushes the hydroxide ion toward the bulk at ambient conditions, despite a clear local free energy minimum in the first contact layer, supporting its preference for optimal water interactions. Importantly, these findings were made possible by leveraging machine learning-based MD simulations of systems encompassing up to 700 atoms and spanning over 200 ns while retaining first-principles accuracy, far beyond the limits set by AIMD. This approach is crucial to our study as it allows us to achieve DFT-level accuracy and extensively sample the phase space, providing insights into the microscopic mechanisms driving these interactions.
@@
The mechanistic insights provided in this work are expected to significantly influence and guide future experimental studies. Previous work [39] has rationalized the macroscopic experimental behavior of nanoconfined electrolytes by postulating the presence of a negative surface charge due to the adsorption of hydroxide ions at the carbon interface. However, our findings not only demonstrate a lack of preferential hydroxide adsorption but also a complex relationship between ion adsorption and surface charge, wherein an adsorbed hydroxide can negatively polarize the surface. The phenomena driving this behavior —hydrogen bonding and ion orientational effects— cannot be captured by continuum theories like Poisson-Boltzmann which are conventionally applied to model these systems. Additionally, the ambiguity in defining interfacial depth often results in conflicting outcomes, compromising the accurate determination of surface acidity [14]. This underscores the necessity of integrating our simulation findings with experiments that offer an atomistic resolution of interfacial signatures [66]. Surface-sensitive spectroscopy techniques, such as second harmonic generation and sum frequency generation spectroscopies[67, 21, 23, 68], provide such resolution as they are active only in noncentrosymmetric regions. Complementary pH measurements within porous carbon could further enhance our understanding of these phenomena. When coupled with nuclear magnetic resonance spectroscopy, these techniques could offer new insights into the behavior of these ions at the interface [69]. Finally, direct atomic resolution imaging via atomic force microscopy could help elucidate differences in water structure [70], providing insights into how interactions affect conductivity in graphene.
@@
This study provides a detailed picture of protonic defects at the graphene-water interface as a foundation for interpreting experimental data and advancing our fundamental understanding of ions at interfaces. Our prediction of high proton propensity to the graphene-water interface opens up the possibility for technological innovations in nanofluidics, heterogeneous solid-liquid catalysis, and other critical domains that rely on proton-mediated processes. Finally, given its model system character, our prediction of an acidic interface between graphene and water is postulated to be transferable to other systems like water in biological channels, geological formations, and technological nano-devices.
@@ Methods
Machine Learning Potentials. In this work, we use the MACE architecture [52], which allows for fast and highly data-efficient training with high-order equivariant message passing and has been proven robust in a wide variety of scenarios [71]. We developed and validated a MACE MLP model (see Section S2) with two layers, a 6 Å cutoff distance, 128 equivariant messages, and a maximal message equivariance of $L=1$. The MLP includes semi-local interactions with a receptive field effectively spanning 12 Å, expected to capture most electrostatic interactions adequately. The final energy and force validation root-mean-square error were 0.7 meV/atom and 17.2 meV/Å, respectively.
@@
To accurately represent the potential energy surface of the systems, we train our MLP model using energies and atomic forces obtained from DFT calculations using the CP2K/Quickstep code [72]. We specifically used the revPBE-D3 [53, 54] functional due to its robust performance in reproducing the structure and dynamics of liquid water [55, 56, 57], while also effectively capturing protonic defects [59] and the interaction energies between water and graphene [58]. Atomic cores are represented using dual-space GTH pseudopotentials [73]. The Kohn-Sham orbitals of oxygen and hydrogen atoms are expanded using the TZV2P basis set, while those of carbon atoms are expanded using the DZVP basis set. An auxiliary plane-wave basis with a cutoff of 1200 Ry was used to represent the density. See Section S1 for further details[74].
@@
The MLP model was systematically developed over five generations. The first generation involved a training set obtained from previous work [36], enhanced by an active learning procedure [75, 46] to incorporate structures that explicitly account for water-carbon interactions at slit widths of 5 and 6.5 Å. This included conditions ranging from low- to high-density water at various temperatures, including 100, 300, and 600 K. The second generation incorporated structures obtained from path integral MD simulations to capture the quantum fluctuations of the nuclei in our model. For the third generation, we targeted various slit widths, including 6, 10, 15, and 20 Å, and included structures corresponding to these dimensions. In the fourth generation, we conducted an additional round of active learning to refine the model based on the conditions sampled thus far. This led to the fifth generation, which included configurations of neutral frames containing a protonic defect pair (both a hydronium and a hydroxide ion) under both bulk water conditions and confined conditions at slit widths of 6, 10, 13 Å. To avoid the complications of applying a background charge for charge neutrality, we did not include isolated protonic defects as suggested in Ref. 59. The final model consisted of 3,378 structures, with 1,303 involving graphene interfaces and 2,075 associated with bulk conditions.
@@
Molecular Dynamics Simulations. All MD simulations reported herein, which were based on the MLP, were performed using the ASE software [76] at a temperature of 300 K, unless explicitly stated otherwise, in the NVT ensemble. A time step of 0.5 fs was employed, and simulations utilized a Langevin thermostat with a friction coefficient of 2.5 ps-1. For each of the five slit widths and two species, we conducted five independent simulations. Each simulation included a 90 ps equilibration period followed by a 4 ns production run, resulting in a total of over 200 ns of simulation time. Uncertainties in reported values were calculated using the standard deviation from these replicates. The simulation cells were initially set up by randomly packing several molecules between the graphene sheets to form one to five well-defined layers of water. To achieve equilibrium density, the graphene sheets were treated in a fully flexible manner, allowing them to adapt without additional constraints. To validate the findings reported, we conducted additional MLP-based biased simulations using umbrella sampling to compare the free energy profiles reported. See Section S1 and S4 for further details.
@@ Acknowledgements.
We thank Marie-Laure Bocquet for fruitful discussions. X.R.A. and A.M. acknowledge support from the European Union under the “n-AQUA” European Research Council project (Grant No. 101071937). K.D.F. acknowledges support from Schmidt Science Fellows and Trinity College, Cambridge. C.S acknowledges partial financial support from the Alexander von Humboldt Stiftung and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) project number 500244608. Via our membership of the UK’s HEC Materials Chemistry Consortium, which is funded by EPSRC (EP/F067496), this work used the ARCHER2 UK National Supercomputing Service (http://www.archer2.ac.uk). This work was also performed using resources provided by the Cambridge Service for Data Driven Discovery (CSD3) operated by the University of Cambridge Research Computing Service (www.csd3.cam.ac.uk), provided by Dell EMC and Intel using Tier-2 funding from the Engineering and Physical Sciences Research Council (capital grant EP/T022159/1), and DiRAC funding from the Science and Technology Facilities Council (www.dirac.ac.uk). Access to CSD3 was also obtained through a University of Cambridge EPSRC Core Equipment Award (EP/X034712/1). We additionally acknowledge computational support and resources from the Cirrus UK National Tier-2 HPC Service at EPCC (http://www.cirrus.ac.uk) funded by the University of Edinburgh and EPSRC (EP/P020267/1).
@@
All data required to reproduce the findings of this work will be made openly available on GitHub upon acceptance of this manuscript.

### v2

@@ Conclusion
Our findings demonstrate a clear preference for the hydronium ion at the graphene-water interface under various confinement regimes, rendering the graphene-water interface enriched with hydronium ions. Enthalpic forces drive the hydronium ion to the interface, maintaining the hydrogen bonding network and enabling energetically favorable interactions with graphene. Conversely, a subtle balance of entropic and enthalpic contributions generally pushes the hydroxide ion toward the bulk at ambient conditions, despite a clear local free energy minimum in the first contact layer, supporting its preference for optimal water interactions. In summary, our work emphasizes the complex interplay of hydrogen bonding, orientational preferences, and thermodynamic forces that dictate the stability of hydronium and hydroxide ions at the graphene-water interface. Coupled with surface polarization effects, leading to counterintuitive charge rearrangements within graphene, these findings provide a comprehensive picture of the two species under confinement and at the graphene-water interface. Importantly, these findings were made possible by leveraging machine learning-based MD simulations of systems encompassing up to 700 atoms and spanning over 200 ns while retaining first-principles accuracy, far beyond the limits set by AIMD. This approach is crucial to our study as it allows us to achieve DFT-level accuracy and extensively sample the phase space, providing insights into the microscopic mechanisms driving these interactions.
@@
The mechanistic insights provided in this work are expected to significantly influence and guide future experimental studies. Previous work [39] has rationalized the macroscopic experimental behavior of nanoconfined electrolytes by postulating the presence of a negative surface charge due to the adsorption of hydroxide ions at the carbon interface. However, our findings not only demonstrate a lack of preferential hydroxide adsorption but also a complex relationship between ion adsorption and surface charge, wherein an adsorbed hydroxide can negatively polarize the surface. The phenomena driving this behavior —hydrogen bonding and ion orientational effects— cannot be captured by continuum theories like Poisson-Boltzmann which are conventionally applied to model these systems. Additionally, the ambiguity in defining interfacial depth can lead to conflicting outcomes, complicating accurate assessments of whether hydronium or hydroxide ions prefer the interface [14]. This underscores the necessity of integrating our simulation findings with experiments that offer an atomistic resolution of interfacial signatures [69]. Surface-sensitive spectroscopy techniques, such as second harmonic generation and sum frequency generation spectroscopies[70, 21, 23, 71], provide such resolution as they are active only in noncentrosymmetric regions. Complementary pH measurements within porous carbon could further enhance our understanding of these phenomena. When coupled with nuclear magnetic resonance spectroscopy, these techniques could offer new insights into the behavior of these ions at the interface [72]. Finally, direct atomic resolution imaging via atomic force microscopy could help elucidate differences in water structure [73], providing insights into how interactions affect conductivity in graphene.
@@
This study provides a detailed picture of protonic defects at the graphene-water interface as a foundation for interpreting experimental data and advancing our fundamental understanding of ions at interfaces. Our prediction of high proton propensity to the graphene-water interface opens up the possibility for technological innovations in nanofluidics, heterogeneous solid-liquid catalysis, and other critical domains that rely on proton-mediated processes. Finally, given the model system character of our setup, we anticipate that this proton-enriched interface behavior may be transferable to other systems like water in biological channels, geological formations, and technological nano-devices.
@@ Methods
Machine Learning Potentials. In this work, we use the MACE architecture [53], which allows for fast and highly data-efficient training with high-order equivariant message passing and has been proven robust in a wide variety of scenarios [74]. We developed and validated a MACE MLP model (see Section S2) with two layers, a 6 Å cutoff distance, 128 equivariant messages, and a maximal message equivariance of $L=1$. The MLP captures semi-local interactions through a receptive field that spans the product of the number of layers and the cutoff distance per layer. In this case, the total receptive field is 12  Å, allowing the MLP to account for interactions within this range. While the model does not explicitly account for long-range effects, the 12 Å receptive field spans nearly the entire width of the slit in most cases, effectively capturing the relevant electrostatic interactions within the simulation. The final energy and force validation root-mean-square error were 0.7 meV/atom and 17.2 meV/Å, respectively.
@@
To accurately represent the potential energy surface of the systems, we train our MLP model using energies and atomic forces obtained from DFT calculations using the CP2K/Quickstep code [75]. We specifically used the revPBE-D3 [54, 55] functional due to its robust performance in reproducing the structure and dynamics of liquid water [56, 57, 58], while also effectively capturing protonic defects [60] and the interaction energies between water and graphene [59]. Atomic cores are represented using dual-space GTH pseudopotentials [76]. The Kohn-Sham orbitals of oxygen and hydrogen atoms are expanded using the TZV2P basis set, while those of carbon atoms are expanded using the DZVP basis set. An auxiliary plane-wave basis with a cutoff of 1200 Ry was used to represent the density. We have examined the influence of the functional and dispersion correction on our results and found that the observed trends remain consistent, irrespective of these choices (see Section S2). For interfacial systems, we used a vacuum of 15 Å to uncouple periodic images in the $z$ direction, leading to negligible interactions between the images as confirmed by the convergence of the energy with respect to the vacuum size. See Section S1 for further details[77].
@@
The MLP model was systematically developed over five generations. The first generation involved a training set obtained from previous work [36], enhanced by an active learning procedure [78, 47] to incorporate structures that explicitly account for water-carbon interactions at slit widths of 5 and 6.5 Å. This included conditions ranging from low- to high-density water at various temperatures, including 100, 300, and 600 K. The second generation incorporated structures obtained from path integral MD simulations to capture the quantum fluctuations of the nuclei in our model. For the third generation, we targeted various slit widths, including 6, 10, 15, and 20 Å, and included structures corresponding to these dimensions. In the fourth generation, we conducted an additional round of active learning to refine the model based on the conditions sampled thus far. This led to the fifth generation, which included configurations of neutral frames containing a protonic defect pair (both a hydronium and a hydroxide ion) under both bulk water conditions and confined conditions at slit widths of 6, 10, 13 Å. To avoid the complications of applying a background charge for charge neutrality, we did not include isolated protonic defects as suggested in Ref. 60. The final model consisted of 3,378 structures, with 1,303 involving graphene interfaces and 2,075 associated with bulk conditions.
@@
Molecular Dynamics Simulations. All MD simulations reported herein, which were based on the MLP, were performed using the ASE software [79] at a temperature of 300 K, unless explicitly stated otherwise, in the NVT ensemble. A time step of 0.5 fs was employed, and simulations utilized a Langevin thermostat with a friction coefficient of 2.5 ps-1. For each of the five slit widths and two species, we conducted five independent simulations. Each simulation included a 90 ps equilibration period followed by a 4 ns production run, resulting in a total of over 200 ns of simulation time. Uncertainties in reported values were calculated using the standard deviation from these replicates. All systems were simulated in orthorhombic simulation cells employing periodic boundary conditions in all three directions. The simulation cells were initially set up by randomly packing several molecules between the graphene sheets to form one to five well-defined layers of water. To prevent interactions between the periodic images, 15 Å  vacuum (exceeding the model’s receptive field) was added in the $z$ direction of these initial configurations. To achieve equilibrium density, the graphene sheets were treated in a fully flexible manner, allowing them to adapt without additional constraints. To validate the findings reported, we conducted additional MLP-based biased simulations using umbrella sampling to compare the free energy profiles reported. See Section S1 and S4 for further details.
@@ Acknowledgements.
We thank Marie-Laure Bocquet for fruitful discussions and Benjamin X. Shi for technical help with the hybrid functional setup. X.R.A. and A.M. acknowledge support from the European Union under the “n-AQUA” European Research Council project (Grant No. 101071937). K.D.F. acknowledges support from Schmidt Science Fellows and Trinity College, Cambridge. C.S acknowledges partial financial support from the Alexander von Humboldt Stiftung and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) project number 500244608. Via our membership of the UK’s HEC Materials Chemistry Consortium, which is funded by EPSRC (EP/F067496), this work used the ARCHER2 UK National Supercomputing Service (http://www.archer2.ac.uk). This work was also performed using resources provided by the Cambridge Service for Data Driven Discovery (CSD3) operated by the University of Cambridge Research Computing Service (www.csd3.cam.ac.uk), provided by Dell EMC and Intel using Tier-2 funding from the Engineering and Physical Sciences Research Council (capital grant EP/T022159/1), and DiRAC funding from the Science and Technology Facilities Council (www.dirac.ac.uk). Access to CSD3 was also obtained through a University of Cambridge EPSRC Core Equipment Award (EP/X034712/1). We additionally acknowledge computational support and resources from the Cirrus UK National Tier-2 HPC Service at EPCC (http://www.cirrus.ac.uk) funded by the University of Edinburgh and EPSRC (EP/P020267/1).
@@
All data required to reproduce the findings of this work are openly available on GitHub (https://github.com/water-ice-group/graphene-water-protons).
Code Availability
All analysis code supporting the findings of this work will be made openly available on GitHub upon acceptance of this manuscript. All simulation code used in this study is available open-source, see CP2K, ASE and MACE for details.